OpenStudy (anonymous):

The sum of two numbers is 44, and the larger number is 2 more than the smaller number. What is the smaller number? If x = the smaller number and y = the larger number, then which of the following systems of equations represents the word problem?

OpenStudy (anonymous):

The sum of two numbers is 44, x = the smaller number and y = the larger number so X+Y=44 x = the smaller number and y = the larger number larger number is 2 more than the smaller number so Y-X=2 The set of equations is: X+Y=44 Y-X=2 After calculations: X=21, Y=23